# Learning to Steer Markovian Agents under Model Uncertainty

### v1

@@ 6 Experiments
In this section, we discuss our experimental results. For more details of all experiments in this section (e.g. experiment setup and training details), we defer to Appx. H. The steering horizon is set to be $T=500$, and all the error bar shows 95% confidence level. We denote $[x]^{+}:=\max\{0,x\}$.
@@ 6.1 Learning Steering Strategies with Knowledge of $f^{*}$
Normal-Form Stag Hunt Game In Fig. 1-(b), we compare the agents’ dynamics with/without steering, where the agents learn to play the Stag Hunt Game in Fig. 1-(a). We report the experiment setup here. Both agents follow the exact NPG (Def. 4.1 with $\widehat{A}^{\bm{\pi}}=A^{\bm{\pi}}$) with fixed learning rate $\alpha=0.01$. For the steering setup, we choose the total utility as $\eta^{\text{goal}}$, and use PPO to train the steering strategy (one can choose other RL or control algorithms besides PPO). We also conduct experiments in a representative zero-sum game ‘Matching Pennies’, which we defer the details to Appx. H.2.
@@
This motivates us to investigate the possibility of steering agents with observations on agents’ behavior only (e.g. trajectories of agents in a game $G$), instead of the full observation of ${\bm{\pi}}$.
In Appx. G, we justify this setup and formalize it as a partially observable extension of our current framework. We consider the evaluation in a grid-world version of the Stag Hunt Game as shown in Fig. 2-(a). In this setting, the state space in game $G$ becomes pixel-based images, and both agents ( blue and red) will adopt Convolutional Neural Networks (CNN) based policies with thousands of parameters and update with PPO. We train a steering strategy, which only takes the agents’ recent trajectories as input to infer the steering reward. As shown in Fig. 2-(b), without direct usage of the agents’ policy, we can still train a steering strategy towards desired solution.
@@ 6.2 Learning Steering Strategies without Knowledge of $f^{*}$
Firstly, we demonstrate the suboptimal behavior if the mediator ignores the model uncertainty and just randomly deploys the optimal strategy of $f_{0.7}$ or $f_{1.0}$. To do this, we train the (history-independent) optimal steering strategy by Alg. 3, as if we know $f^{*}=f_{0.7}$ (or $f^{*}=f_{1.0}$), which we denote as $\psi^{*}_{0.7}$ (or $\psi^{*}_{1.0}$). To meet with our Desideratum 1, we first set the accuracy level $\varepsilon=0.01$, and search the minimal $\beta$ so that the learned steering strategy can achieve $\varepsilon$-steering gap (see Appx. H.3.1). Because of the difference in $\mu$, we have $\beta=70$ and $\beta=20$ in training $\psi^{*}_{0.7}$ and $\psi^{*}_{1.0}$, respectively, and empirically, we observe that $\psi^{*}_{0.7}$ requires much larger steering reward than $\psi^{*}_{1.0}$. As we marked in red in Table 2-(a) and (b), because of the difference in the steering signal, $\psi^{*}_{0.7}$ consumes much higher steering cost to achieve the same accuracy level in $f_{1.0}$, and $\psi^{*}_{1.0}$ may fail to steer agents with $f_{0.7}$ to the desired accuracy. Next, we train another strategy $\psi^{*}_{\text{Belief}}$ via Alg. 4, which predicts the steering reward based on both the agents’ policy ${\bm{\pi}}$ and the belief state of the model. As we can see, $\psi^{*}_{\text{Belief}}$ can almost always achieve the desired $\varepsilon$-steering gap with reasonable steering cost.
Table 2: Evaluation for Proc. 1 (Averaged over 25 different initial ${\bm{\pi}}_{1}$, see Appx. H.1).
(a) Performance in $f_{0.7}$
@@
Large Model Set $|\mathcal{F}|$: The FETE Framework In this part, we evaluate the FETE framework (Proc. 2 in Sec. 5.2). We consider $N$-player normal-form cooperative games with two actions A and B. Each agent will receive the same reward if all of them take the same actions, and no reward otherwise. The rewards for action A and B are 2 and 1, respectively.
@@
Figure 3: Evaluation for Proc. 2. Left: Probability of identifying $f^{*}$ by $f_{\text{MLE}}$. Ours can achieve near 100% success rate with 30 steering steps, while the random exploration takes more than 300 steps. Middle and Right: Comparison between FETE and Oracles. FETE can achieve competitive performance in both steering gap and cost.
In this experiment, we consider “opportunistic” agents who tend to increase the learning rates if the advantage of one action over the other is clear. More concretely, the learning rate of agent $n$ is $\alpha_{n}=[\xi_{n}]^{+}$ with $\xi_{n}\sim\mathcal{N}(1.0+[|Q^{n,{\bm{\pi}}}_{|{\bm{r}}+{\bm{u}}}(\texttt{A})-Q^{n,{\bm{\pi}}}_{|{\bm{r}}+{\bm{u}}}(\texttt{B})|-\lambda^{n}]^{+},0.5^{2})$, where $\lambda^{n}$ is the unknown threshold parameter. In our experiments, we set $N=10$, and for each $n\in[N]$, $\lambda^{n}\in\{0.5,1.0,1.5,+\infty\}$, which results in an extremely large model class $\mathcal{F}$ with $|\mathcal{F}|=4^{10}$. In each step $t\in[T]$, the mediator can observe one learning rates sample $\{\alpha_{n}\}_{n\in[N]}$ of agents, and we need to estimate the true $\{\lambda^{n}\}_{n\in[N]}$ from those samples. By design, the true model can be quickly identified if the steering reward $u$ can lead to a large value gap $|Q^{n,{\bm{\pi}}}_{|{\bm{r}}+{\bm{u}}}(\texttt{A})-Q^{n,{\bm{\pi}}}_{|{\bm{r}}+{\bm{u}}}(\texttt{B})|$. We evaluate our Proc. 2 and report the results in Fig 3. We consider the fixed initial policy with $\forall n\in[N],~{}\pi^{n}_{1}(\texttt{A})=1-\pi^{n}_{1}(\texttt{B})=1/3$. For the exploration, comparing with the noise-based random exploration (Canyakmaz et al.,, 2024), we can see the clear advantage of our strategy trained by reward signals reflecting exploration efficacy [^6] Empirically, we use posterior probability as training signal, see Appx. H.3.2 for more explanation.. Besides, we also compare the steering gap and cost between our FETE method and the Oracle steering strategy – if the mediator knows $f^{*}$ in advance (by solving Obj. (1) with =$\mathcal{F}=\{f^{*}\}$). We choose exploration horizon $\widetilde{T}=30$ suggested by the previous exploration experiment, and report four realizations of $f^{*}\in\{f_{A},f_{B},f_{C},f_{D}\}$. For $f_{A},f_{B}$ and $f_{C}$, the agents share $\lambda^{n}=0.5,1.5,+\infty$, respectively. $f_{D}$ has a diverse setup where $\lambda^{n}=0.5$ for $1\leq n\leq 5$ and $\lambda^{n}=1.5$ for $5<n\leq 10$. As we can see, the steering gap of our FETE remains low comparing with the initial gap $\max_{\bm{\pi}}\eta^{\text{goal}}({\bm{\pi}})-\eta^{\text{goal}}({\bm{\pi}}_{1})\approx 2.0$, and the steering cost is competitive.
@@ 7 Conclusion
In this paper, we introduce the problem of steering Markovian agents to desired game outcomes under model uncertainty. We provide theoretical foundations for this problem by formulating a novel optimization objective and providing existence results. Moreover, we design several algorithmic approaches suitable for varying degrees of model uncertainty in this problem class. We test their performances in different experimental settings and show their effectiveness. Our work opens up avenues for compelling open problems that merit future investigation. Firstly, future work could aim to identify superior optimization objectives that guarantee strictly better performances in terms of steering gap and cost than ours. Secondly, when applying our strategies in real-world applications, constraints on the steering reward budget could be added. Finally, the framework could be generalized to permit non-Markovian agents.
Acknowledgments and Disclosure of Funding
The work is supported by ETH research grant and Swiss National Science Foundation (SNSF) Project Funding No. 200021-207343 and SNSF Starting Grant. VT is supported by an ETH AI Center Doctoral Fellowship. HHN is supported by the SNSF Eccellenza Grant ‘Markets and Norms’.
Appendix A Discussion on Limitations and Societal Impact
Discussion on Limitations
Discussion on Societal Impact
Appendix B Frequently Used Notations
@@
Appendix C Missing Details in the Main Text
C.1 A Real-World Scenario that Can be Modeled as a Stag Hunt Game
@@
C.2 Additional Related Works
@@
MPEC generalises bilevel optimization to problems where the lower level consists of solving an equilibrium problem (Luo et al.,, 1996). (Li et al.,, 2020; Liu et al.,, 2022; Wang et al.,, 2022, 2023; Yang et al.,, 2022). These works consider variants of an MPEC and present gradient based approaches, most of which rely on computing hypergradients via the implicit function theorem and thus strong assumptions on the lower level problem, such as uniqueness of the equilibrium. Most games fail to satisfy such constraints. In contrast, our work makes no assumptions on the equilibrium structure and instead mild assumptions on the learning dynamics.
@@
C.3 A Brief Introduction to Markov Decision Process
@@
C.4 Algorithm for Learning Optimal (History-Independent) Strategy when $f^{*}$ is Known
@@ 4       Sample trajectories with $\psi_{\zeta_{k}}$, $\forall t\in[T]$:
C.5 Algorithm for Learning Belief-State Dependent Steering Strategy
@@ 4       Sample trajectories with $\psi^{k}$ from simulator of $f$:
Appendix D Missing Proofs in Section 3
@@
Appendix E Missing Proofs for Existence when the True Model $f^{*}$ is Known
In this section, we study the Policy Mirror Descent as a concrete example. In Appx. E.1, we provide more details about PMD. Then, we study the PMD with exact updates and stochastic updates in Appx. E.2.1 and E.2.2, respectively. The theorems in Sec. 4.1 will be subsumed as special cases.
E.1 More Details about Policy Mirror Descent
Definition E.1 (Policy Mirror Descent).
@@ E.1 More Details about Policy Mirror Descent
By Pythagorean Theorem and the strictly convexity of $D_{\phi^{n}}$, the projection ${\bm{\pi}}$ in Def. E.1 is unique.
Lemma E.2.
@@
Example E.3 (Natural Policy Gradient).
@@
Example E.4 (Online Gradient Ascent (Zinkevich,, 2003)).
@@
In the following Lem. E.6, we show that constant shift in $\theta^{n}_{t,h}(\cdot|s_{h})$ does not change the projection result. Therefore, when we say the dual variable $\theta$ associated with a given policy ${\bm{\pi}}$, we only consider those $\theta$ satisfying $\mathbb{E}_{a_{h}^{n}\sim\pi^{n}_{h}}[\theta^{n}_{h}(a_{h}^{n}|s_{h})]=0$.
Lemma E.6 (Constant Shift does not Change the Projection).
@@
E.2 Proofs for the Existence of Desired Steering Strategy
@@
In the following, in Appx. E.2.1, as a warm-up, we start with the exact case when the estimation ${\widehat{A}}^{\bm{\pi}}$ is exactly the true advantage value $A^{\bm{\pi}}$ (which can be regarded as a special case of Assump. B). Then, in Appx. E.2.2, we study the general setting and prove Thm. 4.2 as a special case of PMD.
E.2.1 Special Case: PMD with Exact Advantage-Value
Lemma E.7 (Existence of Steering Path between Feasible Policies).
Consider two feasible policies ${\bm{\pi}},{\widetilde{\bm{\pi}}}$ which are induced by dual variables $\{\theta^{n}_{1,h}\}_{h\in[H],n\in[N]}$ and $\{\widetilde{\theta}_{h}^{n}\}_{h\in[H],n\in[N]}$, respectively. If the agents follow Def. E.1 with exact Q value and start with ${\bm{\pi}}_{1}={\bm{\pi}}$, as long as $U_{\max}\geq 2H+\frac{2}{\alpha T}(\max_{n,h,s_{h},a^{n}_{h}}|\widetilde{\theta}_{h}^{n}(a^{n}_{h}|s_{h})-\theta_{h}^{n}(a^{n}_{h}|s_{h})-\mathbb{E}_{\bar{a}_{h}^{n}\sim\pi^{n}_{t,h}(\cdot|s_{h}^{n})}[\widetilde{\theta}_{h}^{n}(\bar{a}^{n}_{h}|s_{h})-\theta_{h}^{n}(\widetilde{a}^{n}_{h}|s_{h})]|)$, there exists a (history-independent) steering strategy $\psi:=\{\psi_{t}\}_{t\in[T]}$ with $\psi_{t}:\Pi^{+}\rightarrow\mathcal{U}$, s.t., ${\bm{\pi}}_{T+1}=\widetilde{\bm{\pi}}$.

### v2

@@ 6 Experiments
In this section, we discuss our experimental results. For more details of all experiments in this section (e.g. experiment setup and training details), we defer to Appx. G. The steering horizon is set to be $T=500$, and all the error bar shows 95% confidence level. We denote $[x]^{+}:=\max\{0,x\}$.
@@ 6.1 Learning Steering Strategies with Knowledge of $f^{*}$
Normal-Form Stag Hunt Game In Fig. 1-(b), we compare the agents’ dynamics with/without steering, where the agents learn to play the Stag Hunt Game in Fig. 1-(a). We report the experiment setup here. Both agents follow the exact NPG (Def. 4.1 with $\widehat{A}^{\bm{\pi}}=A^{\bm{\pi}}$) with fixed learning rate $\alpha=0.01$. For the steering setup, we choose the total utility as $\eta^{\text{goal}}$, and use PPO to train the steering strategy (one can choose other RL or control algorithms besides PPO). We also conduct experiments in a representative zero-sum game ‘Matching Pennies’, which we defer the details to Appx. G.2.
@@
This motivates us to investigate the possibility of steering agents with observations on agents’ behavior only (e.g. trajectories of agents in a game $G$), instead of the full observation of ${\bm{\pi}}$. In Appx. F, we justify this setup and formalize it as a partially observable extension of our current framework. We consider the evaluation in a grid-world version of the Stag Hunt Game as shown in Fig. 2-(a). In this setting, the state space in game $G$ becomes pixel-based images, and both agents ( blue and red) will adopt Convolutional Neural Networks (CNN) based policies with thousands of parameters and update with PPO. We train a steering strategy, which only takes the agents’ recent trajectories as input to infer the steering reward. As shown in Fig. 2-(b), without direct usage of the agents’ policy, we can still train a steering strategy towards desired solution.
@@ 6.2 Learning Steering Strategies without Knowledge of $f^{*}$
Table 1: Evaluation for Proc. 1 (Averaged over 25 different initial ${\bm{\pi}}_{1}$, see Appx. G.1).
(b) Performance in $f_{0.7}$
@@
Firstly, we demonstrate the suboptimal behavior if the mediator ignores the model uncertainty and just randomly deploys the optimal strategy of $f_{0.7}$ or $f_{1.0}$. To do this, we train the (history-independent) optimal steering strategy by Alg. 3, as if we know $f^{*}=f_{0.7}$ (or $f^{*}=f_{1.0}$), which we denote as $\psi^{*}_{0.7}$ (or $\psi^{*}_{1.0}$). To meet with our Desideratum 1, we first set the accuracy level $\varepsilon=0.01$, and search the minimal $\beta$ so that the learned steering strategy can achieve $\varepsilon$-steering gap (see Appx. G.3.1). Because of the difference in $\mu$, we have $\beta=70$ and $\beta=20$ in training $\psi^{*}_{0.7}$ and $\psi^{*}_{1.0}$, respectively, and empirically, we observe that $\psi^{*}_{0.7}$ requires much larger steering reward than $\psi^{*}_{1.0}$. As we marked in red in Table 1-(a) and (b), because of the difference in the steering signal, $\psi^{*}_{0.7}$ consumes much higher steering cost to achieve the same accuracy level in $f_{1.0}$, and $\psi^{*}_{1.0}$ may fail to steer agents with $f_{0.7}$ to the desired accuracy. Next, we train another strategy $\psi^{*}_{\text{Belief}}$ via Alg. 4, which predicts the steering reward based on both the agents’ policy ${\bm{\pi}}$ and the belief state of the model. As we can see, $\psi^{*}_{\text{Belief}}$ can almost always achieve the desired $\varepsilon$-steering gap with reasonable steering cost.
Large Model Set $|\mathcal{F}|$: The FETE Framework In this part, we evaluate the FETE framework (Proc. 2 in Sec. 5.2). We consider an cooperative setting with $N=10$ players. Each agent has two actions A and B, and the mediator only receives non-zero utility when all the agents cooperate together to take action A, i.e. $\eta^{\text{goal}}({\bm{\pi}}):=\prod_{n=1}^{N}\pi^{n}(\texttt{A})$. The agents do not have intrinsic rewards (${\bm{r}}=0$), but the mediator’s can steer them to maximize its own utility by providing additional steering rewards.
@@
Figure 3: Evaluation for Proc. 2. Left: Accuracy of MLE estimator ($\lambda^{n}_{\text{MLE}}$) after doing exploration for $t$ steps. Ours can achieve near 100% accuracy after 30 steering steps, while the random exploration takes more than 300 steps. Middle and Right: Average steering gap and steering cost of Oracle, FETE and FETE-RE. Our FETE achieves competitive performance comparing with Oracle, and significantly outperforms FETE-RE (adaption of SIAR-MPC (Canyakmaz et al.,, 2024) to our setting) in terms of steering gap.
We consider “avaricious agents” with varying degrees of greediness, who tend to decrease the learning rates if the payments by mediator are high. Consequently, they require more steering steps to converge to the desired policies, potentially earning more incentive payments from the mediators. More concretely, the learning rate of agent $n$ is $\alpha_{n}:=[\xi_{n}]^{+}$ with $\xi_{n}\sim\mathcal{N}(1.5-\beta^{n}\cdot[V_{|{\bm{u}}}^{n,{\bm{\pi}}}-\lambda^{n}]^{+},0.5^{2})$, where $\beta^{n}>0$ is a scaling factor and $\lambda^{n}>0$ is the threshold to exhibit avaricious behavior. In our experiments, the model uncertainty comes from multiple possible realization of $\lambda^{n}\in\{0.25,0.75,+\infty\}$, which results in an extremely large model class $\mathcal{F}$ with $|\mathcal{F}|=3^{10}$. Here $\lambda^{n}=+\infty$ corresponds to normal agents whose learning rates are stable. The mediator does not know the agents’ types $\{\lambda^{n}\}_{n\in[N]}$ in advance, and it can only observe one learning rate samples $\{\alpha_{n}\}_{n\in[N]}$ of agents per iteration $t\in[T]$ and estimate the true types from those samples. We consider the fixed initial policy with $\forall n\in[N],~{}\pi^{n}_{1}(\texttt{A})=1-\pi^{n}_{1}(\texttt{B})=1/3$, and set the maximal steering reward $U_{\max}=1.0$.
To understand the exploration challenge, note that, during the exploration phase, if the steering signal ${\bm{u}}$ is not strong enough, i.e. $V^{n,{\bm{\pi}}}_{|{\bm{u}}}<\lambda^{n}$, the mediator may fail to distinguish those avaricious agents from the normal ones, because they behave exactly the same. Such failure can lead to undesired outcomes in the exploitation phase: higher steering rewards can accelerate the convergence of normal agents, but can lead to larger steering gaps for avaricious agents.
We provide the evaluation results in Fig. 3. First, we compare the exploration efficiency. We can see the clear advantage of our strategic exploration in FETE (Procedure 2) compared with noise-based random exploration (Canyakmaz et al.,, 2024). Next, we compare the steering gaps and costs of three methods: (i) FETE; (ii) FETE-RE; (iii) Oracle – if the mediator knows $f^{*}$ in advance and solving Obj. (1) with =$\mathcal{F}=\{f^{*}\}$. Here FETE-RE can be regarded as adaption of SIAR-MPC (Canyakmaz et al.,, 2024) to our case by replacing strategic exploration in FETE with random exploration (see Appx. G.4 for more explanation). We choose exploration horizon $\widetilde{T}=30$ suggested by the previous exploration experiment, and report results for three realizations of $f^{*}\in\{f_{1},f_{2},f_{3}\}$. For $f_{1}$ and $f_{2}$, all the agents share $\lambda^{n}=0.75$ and +$\infty$+ ∞, respectively. $f_{3}$ is a mixed setup where $\lambda^{n}=0.75$ for $1\leq n\leq 5$ and $\lambda^{n}=+\infty$ for $5<n\leq 10$. As we can see, comparing with Oracle, both the steering gap and cost of our FETE are competitive. Moreover, thanks to our strategic exploration method, FETE exhibits significiant advantage over Canyakmaz et al., (2024) in terms of steering gaps.
@@ 7 Conclusion
In this paper, we introduce the problem of steering Markovian agents under model uncertainty. We provide theoretical foundations for this problem by formulating a novel optimization objective and providing existence results. Moreover, we design several algorithmic approaches suitable for varying degrees of model uncertainty in this problem class. We test their performances in different experimental settings and show their effectiveness. Our work opens up avenues for compelling open problems that merit future investigation. Firstly, future work could aim to identify superior optimization objectives that guarantee strictly better performances in terms of steering gap and cost than ours. Secondly, when applying our strategies in real-world applications, constraints on the steering reward budget could be added. Finally, the framework could be generalized to permit non-Markovian agents.
Acknowledgement
The work is supported by ETH research grant and Swiss National Science Foundation (SNSF) Project Funding No. 200021-207343 and SNSF Starting Grant. VT is supported by an ETH AI Center Doctoral Fellowship. HHN is supported by the Swiss National Science Foundation under Eccellenza Grant ‘Markets and Norms,’ as well as under NCCR Automation Grant 51NF40-180545.
Reproducibility Statement
The codes for all the experiments in this paper and the instructions for running can be found in the supplementary materials.
Appendix A Frequently Used Notations
@@
Appendix B Missing Details in the Main Text
B.1 A Real-World Scenario that Can be Modeled as a Stag Hunt Game
@@
B.2 Additional Related Works
We first complements the comparison with Zhang et al., (2023) in Sec. 1.1 by noting a minor but worth to be mentioned difference between our setting and (Zhang et al.,, 2023) in terms of incentive schemes. While they consider that the mediator influences the agents’ learning dynamics through a scalar payment function, we operate with additional steering rewards in a multi-dimensional reward vector space. As a result, there may not exist direct ways to translate the steering strategies between both settings, especially in the bandit feedback setting where only the sampled actions of agents can be observed (Zhang et al.,, 2023).
@@ B.2 Additional Related Works
MPEC generalises bilevel optimization to problems where the lower level consists of solving an equilibrium problem (Luo et al.,, 1996). (Li et al.,, 2020; Liu et al.,, 2022; Wang et al.,, 2022; 2023; Yang et al.,, 2022). These works consider variants of an MPEC and present gradient based approaches, most of which rely on computing hypergradients via the implicit function theorem and thus strong assumptions on the lower level problem, such as uniqueness of the equilibrium. Most games fail to satisfy such constraints. In contrast, our work makes no assumptions on the equilibrium structure and instead mild assumptions on the learning dynamics.
@@
Episodic RL and Non-Episodic RL
B.3 A Brief Introduction to Markov Decision Process
@@
B.4 Algorithm for Learning Optimal (History-Independent) Strategy when $f^{*}$ is Known
@@ 4       Sample trajectories with $\psi_{\zeta_{k}}$, $\forall t\in[T]$:
B.5 Algorithm for Learning Belief-State Dependent Steering Strategy
@@ 4       Sample trajectories with $\psi^{k}$ from simulator of $f$:
Appendix C Missing Proofs in Section 3
@@
Appendix D Missing Proofs for Existence when the True Model $f^{*}$ is Known
In this section, we study the Policy Mirror Descent as a concrete example. In Appx. D.1, we provide more details about PMD. Then, we study the PMD with exact updates and stochastic updates in Appx. D.2.1 and D.2.2, respectively. The theorems in Sec. 4.1 will be subsumed as special cases.
D.1 More Details about Policy Mirror Descent
Definition D.1 (Policy Mirror Descent).
@@ D.1 More Details about Policy Mirror Descent
By Pythagorean Theorem and the strictly convexity of $D_{\phi^{n}}$, the projection ${\bm{\pi}}$ in Def. D.1 is unique.
Lemma D.2.
@@
Example D.3 (Natural Policy Gradient).
@@
Example D.4 (Online Gradient Ascent (Zinkevich,, 2003)).
@@
In the following Lem. D.6, we show that constant shift in $\theta^{n}_{t,h}(\cdot|s_{h})$ does not change the projection result. Therefore, when we say the dual variable $\theta$ associated with a given policy ${\bm{\pi}}$, we only consider those $\theta$ satisfying $\mathbb{E}_{a_{h}^{n}\sim\pi^{n}_{h}}[\theta^{n}_{h}(a_{h}^{n}|s_{h})]=0$.
Lemma D.6 (Constant Shift does not Change the Projection).
@@
D.2 Proofs for the Existence of Desired Steering Strategy
@@
In the following, in Appx. D.2.1, as a warm-up, we start with the exact case when the estimation ${\widehat{A}}^{\bm{\pi}}$ is exactly the true advantage value $A^{\bm{\pi}}$ (which can be regarded as a special case of Assump. B). Then, in Appx. D.2.2, we study the general setting and prove Thm. 4.2 as a special case of PMD.
D.2.1 Special Case: PMD with Exact Advantage-Value
Lemma D.7 (Existence of Steering Path between Feasible Policies).
Consider two feasible policies ${\bm{\pi}},{\widetilde{\bm{\pi}}}$ which are induced by dual variables $\{\theta^{n}_{1,h}\}_{h\in[H],n\in[N]}$ and $\{\widetilde{\theta}_{h}^{n}\}_{h\in[H],n\in[N]}$, respectively. If the agents follow Def. D.1 with exact Q value and start with ${\bm{\pi}}_{1}={\bm{\pi}}$, as long as $U_{\max}\geq 2H+\frac{2}{\alpha T}(\max_{n,h,s_{h},a^{n}_{h}}|\widetilde{\theta}_{h}^{n}(a^{n}_{h}|s_{h})-\theta_{h}^{n}(a^{n}_{h}|s_{h})-\mathbb{E}_{\bar{a}_{h}^{n}\sim\pi^{n}_{t,h}(\cdot|s_{h}^{n})}[\widetilde{\theta}_{h}^{n}(\bar{a}^{n}_{h}|s_{h})-\theta_{h}^{n}(\widetilde{a}^{n}_{h}|s_{h})]|)$, there exists a (history-independent) steering strategy $\psi:=\{\psi_{t}\}_{t\in[T]}$ with $\psi_{t}:\Pi^{+}\rightarrow\mathcal{U}$, s.t., ${\bm{\pi}}_{T+1}=\widetilde{\bm{\pi}}$.